# NestQuant: Nested Lattice Quantization for Matrix Products and LLMs

### v1

@@ 3 Outline of NestQuant approach
The details of choosing $k$ and values of $\beta_{t}$ are described in Section D. Here we only note that the overall compression rate becomes $R+{1\over d}{\log_{2}k}$. In some cases, we are using nvcomp [24] for compressing a vector of $n/8$ chosen betas, in which case rate penalty is reduced below ${1\over d}\log_{2}k$. We note that in our comparisons, including Table 1, we always use this effective rate for a fair comparison with other algorithms.
@@
NestQuant, SpinQuant and theory: As mentioned above, the use of nested lattice codes is rooted in theory. In [8] it was shown that nested lattice quantizers of high-dimensions attain the optimal rate-distortion tradeoff for matrix multiplication. Since the lattices used for proving that result do not admit efficient lattice decoding, here we resort to $n$-dimensional lattices constructed as the Cartesian product of $n/8$ copies of the Gosset lattice, whose dimension is $d=8$. To understand how much loss in efficiency this leads, Fig. 3 compares NestQuant, SpinQuant (uniform quantization with cubic shaping) and information-theoretic lower bound (1). Details of this experiment can be found in Section 5.1. We can see that NestQuant is reasonably close to the fundamental limit and significantly outperforms SpinQuant.
@@ 4.1 Nested lattice codebook
In this section, we describe the construction for a Vector Quantization (VQ) codebook of size $q^{d}$ for quantizing an $d$-dimensional vector, where $q$ is an integer parameter. To quantize a vector, we find the closest codebook element by Euclidean norm. We describe efficient encoding and decoding algorithms to a quantized representation in $\mathbb{Z}_{q}^{d}$.
Let $\Lambda$ be a lattice in $\mathbb{R}^{d}$ with generator matrix $G$. We define the coordinates of a point $x\in\Lambda$ to be an integer vector $v$ such that $x=Gv$. Each point $P\in\Lambda$ has a corresponding Voronoi region $\mathcal{V}_{\Lambda}(P)$, for which $P$ is the closest point in $\Lambda$ with respect to $L^{2}$ metric. To define the codebook, we consider the scaled lattice $q\Lambda$. Then:
Definition 4.1.
$x\in\Lambda$ belongs to codebook $C$ iff $x\in\mathcal{V}_{q\Lambda}(0)$. Let $v$ be the coordinates of $x$. Then, the quantized representation of $x$ is $\mathcal{Q}(x):=v\;\mathrm{mod}\ q$. Note that $\mathcal{Q}$ is a bijection between $C$ and $\mathbb{Z}_{q}^{d}$
Using this representation, we can describe the encoding and decoding functions, assuming the point $x$ we are quantizing is in $\mathcal{V}_{q\Lambda}(0)$. We will also need an oracle $Q_{\Lambda}(x)$, which maps $x$ to the closest point in $\Lambda$ to $x$.
@@
Input: $x\in V_{q\Lambda}(0)$, $Q_{\Lambda}$
@@
In practice, we will be using the Gosset ($E_{8}$) lattice as $\Lambda$ with $d=8$. This lattice is a union of $D_{8}$ and $D_{8}+\frac{1}{2}$, where $D_{8}$ contains elements of $\mathbb{Z}^{8}$ with even sum of coordinates. There is a simple algorithm for finding the closest point in the Gosset lattice, first described in [25]. We provide the pseudocode for this algorithm together with the estimation of its runtime in Appendix B.
@@ 4.2 Matrix quantization
Input: $A$ — a vector of size $n=db$, $q$, array of $\beta$
$QA$$n$ integers $\triangleright$ quantized representation
$B$$b$ integers $\triangleright$ scaling coefficient indices
@@ 4.3 LLM quantization
Recall that we apply a rotation matrix $H$ to every weight-activation pair of a linear layer without changing the output of the network. Let $n$ be the number of input features to the layer.
@@
When quantizing a weight, we modify the NestQuant algorithm by introducing corrections to unquantized weights when a certain vector piece is quantized. We refer the reader to section 4.1 of [17] for a more detailed description.
@@ 4.4 Optimal scaling coefficients
One of the important parts of the algorithm is finding the optimal set of $\beta_{i}$. Given the distribution of 8-vectors that are quantized via a codebook, it is possible to find an optimal set of given size exactly using a dynamic programming approach, which is described in Appendix D.
@@ 4.5 Algorithm summary
    Collect the statistics for LDLQ. For each linear layer with in-dimension $d$, we compute a $d\times d$ matrix $H$.
@@
    We run a dynamic programming algorithm described in Appendix D on the 8-vectors to find the optimal $\beta$-values for each weight matrix.
@@
    We quantize the weights using LDLQ and precomputed $\beta$.
@@
    During inference, we quantize any activation before it’s passed to the linear layer, and any KV cache entry before it is saved.
Note the complete lack of fine-tuning needed to make our method work.
@@ 5.1 Simulated Data
For NestQuant, we do a grid search over $(q,k)$. For given $q$ and $k$, we find the best subset in $\{\frac{m}{2}\text{ for }m=1,2,\ldots,50\}$ of scaling coefficients $\beta$ of size $k$ using the algorithm from Appendix D. Then we calculate the expected bits per entry computed as $\log_{2}{q}+\frac{1}{8}\sum_{i=1}^{k}p(\beta_{i})\log_{2}p(\beta_{i})$ where $p(\beta_{i})$ is the probability that the $i$’th beta is used in quantization. In Figure 3, we plot the efficient frontier of bits per entry vs root mean $L_{2}$ loss.
@@ 5.2 Llama results
We quantize Llama-3-8B model [26] using different values of $q$. We choose the number of scaling coefficients ($k$) to be equal to $4$, the Section 5.5 explains the rationale behind this choice. More details on the hyperparameter choice of the experiments are in Appendix E. For each experiment, we compute the number of bits per entry similar to Section 5.1, but for the setup of compressed $\beta$ indices, we run the zstd compression algorithm instead of using the entropy of the distribution. As our evaluation metric, we use the perplexity on the validation split of wikitext2 with context length $2048$.
We also perform the evaluation of NestQuant on various zero-shot benchmarks: ARC-Easy and ARC-Challenge [27], Hellaswag [28], [29], and Winogrande [30]. The results on 4-bit models with comparisons to other models are summarized in Table 1.
@@
Table 2: Wikitext2 perplexity of NestQuant quantization of Llama-3-8B at different rates. The "bits" column is the bit rate per entry with zstd compression of scaling coefficients, and "bits (no zstd)" is the bit rate without compression. The "W", "W+KV", and "W+KV+A" describe the quantization regime (whether weights, KV cache, or activations are quantized). The perplexity of non-quantized model is $6.139$
@@ 5.3 Comparison to other methods
We compare our method to the current state-of-the-art method SpinQuant. On the WikiText2 dataset, we computed the perplexity scores of the quantized models on context size 2048. Our method demonstrates superior perplexity scores by a high margin. On W4KV4A4 (4-bit weights, KV-cache, and activations) quantization of Llama 3-8B we achieve a perplexity score of 6.6 compared to the reported score of 7.3 in SpinQuant (See table 1). Impressively, our method outperforms SpinQuant, without the need of learned rotations. Even without LDLQ, we achieve a perplexity score of 6.8, which is still better than SpinQuant.
5.4 Results for other models
q Bits Bits (no zstd) W W + KV W + KV + A
14 3.99 4.06 10.061 10.529 11.197
12 3.76 3.837 10.178 10.862 11.910
10 3.50 3.57 10.377 11.552 14.191
8 3.18 3.25 10.850 13.309 18.710
5.5 The choice of $k$
More ablation studies for LDLQ and the choice of rotation are described in Appendix F.
6 Impact statement
@@
7 Acknowledgements
This work was supported in part by the National Science Foundation under Grant No CCF-2131115. We thank Foundry.ai for compute resources.
@@ Appendix A Figures
Figure 6: Complement Gaussian measure of a $8$-dimensional cube (corresponding to shaping using an $\ell_{\infty}$ ball), a Voronoi region of the Gosset lattice $E_{8}$ (corresponding to shaping using Voronoi codes with base lattice $E_{8}$), and a Euclidean ball (corresponding to shaping with a ball, which does not admit efficient implementation)
Appendix B Gosset oracle

### v3

@@ 3 Outline of NestQuant approach
The details of choosing $k$ and values of $\beta_{t}$ are described in Section F. Here we only note that the overall compression rate becomes $R+{\frac{1}{d}}{\log_{2}k}$. In some cases, we are using nvcomp nvcomp for compressing a vector of $n/8$ chosen betas, in which case rate penalty is reduced below ${\frac{1}{d}}\log_{2}k$. We note that in our comparisons, including Table 1, we always use this effective rate for a fair comparison with other algorithms.
@@
NestQuant, SpinQuant and theory: As mentioned above, the use of nested lattice codes is rooted in theory. In (op2024, ) it was shown that nested lattice quantizers of high-dimensions attain the optimal rate-distortion tradeoff for matrix multiplication. Since the lattices used for proving that result do not admit efficient lattice decoding, here we resort to $n$-dimensional lattices constructed as the Cartesian product of $n/d$ copies of the Gosset lattice, whose dimension is $d=8$. To understand how much loss in efficiency this leads to, Fig. 3 compares NestQuant, SpinQuant (uniform quantization with cubic shaping) and information-theoretic lower bound (1), for synthetic data. Details of this experiment can be found in Section 5.1. We can see that NestQuant is reasonably close to the fundamental limit and significantly outperforms SpinQuant.
@@ 4.1 Nested lattice codebook
In this section, we describe the construction for a Vector Quantization (VQ) codebook of size $q^{d}$ for quantizing a $d$-dimensional vector, where $q$ is an integer parameter. This construction is based on Voronoi codes ConwaySloane83 and admits efficient encoding and decoding algorithms, whenever the base lattice has an efficient closest lattice vector algorithm. Another appealing feature of Voronoi codes is that the encoding/decoding complexity is independent of the quantozation rate $R=\log_{2}(q)$.
Definition 4.1 (Voronoi code ConwaySloane83 ).
The Voronoi codebook with base lattice $\Lambda\subset\mathbb{R}^{d}$ and nesting ratio $q\in\mathbb{N}$, corresponding to rate $R=\log_{2}(q)$ bits/entry, is defined as $C=\Lambda\cap\mathcal{V}_{q\Lambda}=\Lambda\cap(q\mathcal{V}_{\Lambda})\subset\mathbb{R}^{d}$. In particular $\lambda\in\Lambda$ belongs to codebook $C$ iff $\lambda\in\mathcal{V}_{q\Lambda}$, and $|C|=q^{d}$.
The Voronoi code consists of the minimum energy representative of each coset in $\Lambda/q\Lambda\cong(\mathbb{Z}/q\mathbb{Z})^{d}$. Consequently, we can represent each coset, and hence, each codeword in $C$, as an element of $\mathbb{Z}_{q}^{d}$ ConwaySloane83 .
Assuming we have access to an oracle $Q_{\Lambda}(x)$, which maps $x\in\mathbb{R}^{d}$ to its closest point in $\Lambda$, the quantization (encoding) and dequantization (decoding) of a Voronoi code are described in Algorithm 1 and Algorithm 2, respectively. Here, $G\in\mathbb{R}^{d\times d}$ is a generator matrix of $\Lambda$. The encoder first maps $x\in\mathbb{R}^{d}$ to its nearest lattice point $p=Q_{\Lambda}(x)$. Since it only has a budget of $dR$ bits for describing $p$, it only describes the coset of $\Lambda/q\Lambda$ it belongs to. This is done by sending $v\;\mathrm{mod}\ q$, where $v\in\mathbb{Z}^{d}$ is the integer vector for which $p=Gv$, referred to as the coordinates of $p$. Upon receiving $v\;\mathrm{mod}\ q$, the decoder knows that $Q_{\Lambda}(x)\in p+q\Lambda$, and has to choose one point in this coset as the reconstruction $\hat{x}\in\mathbb{R}^{d}$. It chooses $\hat{x}$ as the minimum energy vector in $p+q\Lambda$, corresponding to the unique point in $p+q\Lambda\cap\mathcal{V}_{q\Lambda}$. We have that $\hat{x}=Q_{\Lambda}(x)$ iff $Q_{\Lambda}(x)\in\mathcal{V}_{q\Lambda}=q\mathcal{V}_{\Lambda}$. When $Q_{\Lambda}(x)\notin q\mathcal{V}_{\Lambda}$ the quantizer is in overload.
@@
Input: $x\in\mathbb{R}^{d}$, $Q_{\Lambda}$
@@
In our experiments for this paper, we used the Gosset ($E_{8}$) lattice as $\Lambda$ with $d=8$. This lattice is a union of $D_{8}$ and $D_{8}+\frac{1}{2}$, where $D_{8}$ contains elements of $\mathbb{Z}^{8}$ with even sum of coordinates. There is a simple algorithm for finding the closest point in the Gosset lattice, first described in 1056484 . We provide the pseudocode for this algorithm together with the estimation of its runtime in Appendix C.
@@ 4.2 Matrix quantization
Input: $A$ — a vector of size $n=db$, $q$, array of $k$ scaling coefficients $\beta_{1},\ldots,\beta_{k}$
$QA$$n$ integers in $\{0,1\ldots,q-1\}$$\triangleright$ quantized representation
$B$$b$ integers in $\{1,\ldots,k\}$triangleright scaling coefficient indices
@@ 4.3 LLM quantization
Recall that we apply a rotation matrix $H$ to every weight-activation pair of a linear layer without changing the output of the network. Let $n$ be the number of input features to the layer. Following tseng2024 ; liu2024 ; lin2024duquantdistributingoutliersdual :
@@
When quantizing a weight, we modify the NestQuant algorithm by introducing corrections to unquantized weights when a certain vector piece is quantized via the QA-LDLQ mechanism, described in section 4.5. It is based on LDLQ algorithm, described in section 4.1 of tseng2024 .
@@ 4.4 Optimal scaling coefficients
One of the important parts of the algorithm is finding the optimal set of $\beta_{i}$. Given the distribution of $d$-dimensional vectors that are quantized via a Voronoi codebook, it is possible to find an optimal set of given size exactly using a dynamic programming approach, which is described in Appendix F.
4.5 LDLQ with quantized activations (QA-LDLQ)
We note that in the presence of activation quantization, the LDLQ and GPTQ algorithms for weight quantization become significantly suboptimal In fact, using original LDLQ on Llama-3-70B produces $\infty$ perplexity at W4A4 setting due to significant outliers in layer 0., thus necessitating a correction (QA-LDLQ) that we describe here.
Let $W$ be the weight (shape $a\times n$), $X$ be a random original (unquantized) activation vector (shape $n\times 1$), and $Z$ be $X$’s quantization error, which we model as random zero-mean noise (shape $n\times 1$) independent of $X$. Then, if $U$ is the quantized weight, the output quantization error becomes $\delta(U):=WX-U(X+Z)$, so that the loss to be minimized is $\mathbb{E}[\|\delta(U)\|^{2}]$ instead of $\mathbb{E}[\|(W-U)X\|^{2}]$ that LDLQ and GPTQ minimize.
Lemma 4.2.
Suppose $Z$ is independent from $X$, $\mathbb{E}[Z]=0$, and let $H=\mathbb{E}[XX^{\top}]$ and $J=\mathbb{E}[ZZ^{\top}]$. Then for any set $\mathcal{C}_{Q}\subset\mathbb{R}^{a\times n}$
$U^{*}$ $=\operatorname*{\arg\!\min}_{U\in\mathcal{C}_{Q}}\mathbb{E}[\|\delta(U)\|^{2}]$
$=\operatorname*{\arg\!\min}_{U\in\mathcal{C}_{Q}}(\tilde{W}-U)(H+J)(\tilde{W}-U)^{\top},$ (6)
where $\tilde{W}=WH(H+J)^{-1}$.
Thus, QA-LDLQ consists of a) computing $\tilde{W}$; b) running the standard LDLQ but with matrix $\tilde{W}$ as input and Hessian set to $H+J$. More details, practical aspects of QA-LDLQ, as well as proof of the lemma are found in Appendix B.
@@ 4.6 Algorithm summary
    Collect the statistics for LDLQ via calibration data. For each linear layer with in-dimension $d$, we compute a $d\times d$ “Hessian” matrix $H$.
@@
    We run a dynamic programming algorithm described in Appendix F on the 8-vectors to find the optimal $\beta$-values for each weight matrix.
@@
    We quantize the weights using QA-LDLQ and precomputed $\beta$.
@@
    During inference, we quantize any activation before it is passed to the linear layer, and any KV cache entry before it is saved.
Note that we do not undertake any expensive (but surely useful) fine-tuning, such as optimizing rotation matrices or post-quantization training, as in liu2024 and tseng2024 , since our goal is demonstrating the basic primitive, not obtaining the absolute SOTA numbers.
@@ 5.1 Simulated Data
For NestQuant, we do a grid search over $(q,k)$. For given $q$ and $k$, we find the best subset in $\frac{1}{2}\cdot\{1,2,\ldots,50\}$ of scaling coefficients $\beta$ of size $k$ using the algorithm from Appendix F. Then we calculate the expected bits per entry computed as $\log_{2}{q}+\frac{1}{8}\sum_{i=1}^{k}p(\beta_{i})\log_{2}p(\beta_{i})$ where $p(\beta_{i})$ is the probability that the $i$’th beta is used in quantization. In Figure 3, we plot the efficient frontier of bits per entry vs root mean $L_{2}$ loss.
@@ 5.2 Llama results
Bits (W-A-KV) Method Llama-2-7B Llama-2-13B Llama-2-70B Llama-3-8B Llama-3-70B 16-16-16 Floating point 5.47 4.88 3.32 6.14 2.86 4-16-16 QuaRot 5.60 5.00 3.41 - - QuIP# 5.56 4.95 3.38 - - OstQuant 5.64 4.94 3.41 6.53 3.19 NestQuant 5.53 4.93 3.38 6.31 3.14 NestQuantM 5.55 4.95 - 6.35 - 4-16-4 NestQuant 5.57 4.96 3.39 6.37 3.19 NestQuantM 5.59 4.99 - 6.49 - 4-4-16 SpinQuant 5.9 5.2 3.8 7.1 - OstQuant 5.60 5.14 3.57 7.24 3.97 DuQuant 6.08 5.33 3.76 8.06 - 4-4-4 QuaRot 6.10 5.40 3.79 8.16 6.66 SpinQuant 5.9 5.3 3.8 7.3 - OstQuant 5.91 5.25 3.59 7.29 4.01 NestQuant 5.67 5.03 3.49 6.63 3.61 NestQuantM 5.73 5.07 - 6.82 -
Table 2: The wikitext2 perplexity at context window of 2048 for various quantization methods of Llama models.
We quantize the Llama-3-8B model grattafiori2024llama3herdmodels using different values of $q$. We choose the number of scaling coefficients ($k$) to be equal to 444, the Section H.1 explains the rationale behind this choice. More details on the hyperparameter choice of the experiments are in Appendix G. For each experiment, we compute the number of bits per entry similar to Section 5.1, but for the setup of compressed $\beta$ indices, we run the zstd compression algorithm instead of using the entropy of the distribution. As our evaluation metric, we use the perplexity on the validation split of wikitext2 with context length 204820482048.
We also perform the evaluation of NestQuant on various zero-shot benchmarks: ARC-Easy and ARC-Challenge clark2018arc , Hellaswag zellers2019 , bisk2019piqa , and Winogrande sakaguchi2019winogrande . The results on 4-bit models with comparisons to other models are summarized in Table 1.
@@
Table 3: Wikitext2 perplexity of NestQuant quantization of Llama-3-8B at different rates. The "bits" column is the bit rate per entry with zstd compression of scaling coefficients, and "bits (no zstd)" is the bit rate without compression. The "W", "W+KV", and "W+KV+A" describe the quantization regime (whether weights, KV cache, or activations are quantized). The perplexity of non-quantized model is 6.1396.1396.139
@@ 5.3 Comparison to other methods
In comparisons to other methods, we focus on 4-bit setup, choosing $q=14$ and $k=4$. We show the WikiText2 perplexity comparisons for multiple Llama models in table 2, and other benchmark comparisons (for Llama-3-8B) in table 1. The methods that we include in the table are SpinQuant liu2024 , QuIP# tseng2024 , QuaRot ashkboos2024 , DuQuant lin2024duquantdistributingoutliersdual , and OstQuant hu2025ostquantrefininglargelanguage . On the WikiText2 dataset, we computed the perplexity scores of the quantized models on context size 2048.
NestQuant consistently achieves better perplexity metrics across different models for both weight-only regime and full quantization. In fact, for all models we have tested, except Llama-2-7B, NestQuant with W4A4KV4 (4-bit weights, activations, and KV-cache) quantization even outperforms previous works with W4A4KV16. On W4KV4A4 (4-bit weights, KV-cache, and activations) quantization of Llama 3-8B we achieve a perplexity score of 6.6 compared to  7.3 in SpinQuant and OstQuant. Even without LDLQ, we achieve a perplexity score of 6.8, which is still better. Finally, NestQuant outperforms QuIP# for weight-only Llama-2 quantization.
In addition, we evaluate a simpler version of the algorithm that is easier to run in hardware that is called NestQuantM. More detailes on it can be found in section D.
We present ablation studies for our design choices in in Appendix H.
Impact statement
@@
Acknowledgments
We want to thank Nikita Lazarev and Shang Yang for important discussions that helped us to implement an efficient CUDA kernel, as well as Guangxuan Xiao for discussions about the algorithm.
This work was supported in part by the National Science Foundation under Grant No CCF-2131115 and the ISF under Grant 1641/21. We thank Foundry.ai and Lambda Labs for compute resources.
Research was supported, in part, by the United States Air Force Research Laboratory and the United States Air Force Artificial Intelligence Accelerator under Cooperative Agreement Number FA8750-19-2-1000. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the United States Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Figures
Figure 5: Complement Gaussian measure of a 888-dimensional cube (corresponding to shaping using an $\ell_{\infty}$ ball), a Voronoi region of the Gosset lattice $E_{8}$ (corresponding to shaping using Voronoi codes with base lattice $E_{8}$), and a Euclidean ball (corresponding to shaping with a ball, which does not admit efficient implementation)
Appendix B Details of QA-LDLQ
LDLQ details. We briefly mention the details of LDLQ (for more details we refer the reader to [34] or [21]). The main observation is that we want to minimize $\lVert(W-U)X\rVert^{2}$ where $U$ is the quantized weight and $X$ is the matrix of activations. We observe that $\lVert(W-U)X\rVert^{2}=\lVert(W-U)L_{X}\sqrt{D_{X}}\rVert^{2}$ where $L_{X}D_{X}L_{X}^{\top}=H=\mathbb{E}[XX^{\top}]$ and $L_{X}$ is a uni-lower-triangular matrix and $D_{X}$ is diagonal. LDLQ proceeds by optimally quantizing the last column of $U$ (for example using the quantizer for NestQuant). It then recursively removes the feedback $i$, $(W-U)(L_{X}-I)E_{i}$ from the rest of the columns (here $E_{i}$ is the zero matrix with the $ii$ index set to 1). This may be written compactly as $U=\mathcal{Q}(W-(W-U)(L_{X}-I))$ where $\mathcal{Q}$ quantizes each column separately.
Motivation. When quantizing Llama-3-70B with original LDLQ, we have noticed very poor (ppl $\sim 10^{4}$) performance. We have observed that the reason for this is quantization of activations for a small subset of linear layers (4 out of 560), and if they are left in full precision, the perplexity becomes more reasonable (less than 4). We have found them by choosing the layers with the largest amplification ratio — a concept we define next.
Consider a weight $W$ (shape $a\times n$), where $n$ is the number of in-features, and $a$ is the number of out-features. We define amplification of a random vector $X$ ($n\times 1)$ by $W$ as $\alpha(W,X):=\frac{\mathbb{E}[\left\lVert WX\right\rVert_{2}]}{\mathbb{E}[\left\lVert X\right\rVert_{2}]}$. Then, if $X$ is the distribution of input activations, and $Z$ is a random Gaussian vector, we define the amplification ratio for $W$ as $\frac{\alpha(W,Z)}{\alpha(W,X)}$. We note that if a layer has large amplification ratio, its activations are harder to quantize, because the magnitude of quantization noise will be increased much more significantly than the magnitude of the activations themselves. One extreme example of this issue is the value projection of the attention of the first transformer block in Llama-3-70B. This layer has an amplification ratio of 157$\sim$ 157∼ 157, as computed from 101010 wikitext2 sequences of length 204820482048. This makes naive 444-bit quantization of activations nearly impossible, as even small perturbations of the input of the layer are greatly amplified in the output.
We have developed QA-LDLQ to mitigate the issue of large amplification ratio. We model quantization noise as a random Gaussian vector with mean 0 and covariance matrix $J=\varepsilon^{2}I$, where $\varepsilon^{2}$ depends on the quantization rate and the statistics of $X$. For large values of $\varepsilon^{2}$ the modified weight matrix $\tilde{W}=WH(H+\varepsilon^{2}I)^{-1}$ is more robust to random perturbations of the input (thus, decreasing the amplification ratio), but on the other hand, large $\varepsilon^{2}$ results in a larger bias, as expressed in the term $C(W,H,J)$ below. Figure 6 demonstrates this tradeoff for the value projection layer mentioned earlier.
Refer to caption Refer to caption
Figure 6: We run QA-LDLQ for value projection layer of the first transformer block of Llama-3-70B. We try different values of $\varepsilon$ on logarithmic scale from $10^{-5}$ to 111. For each $\varepsilon$, we find modified weight $\tilde{W}$, and plot the amplification ratio for $\tilde{W}$ in $y$-axis, as well as how close the outputs of the weight $\tilde{W}$ to the outputs of weight $W$. The value on $x$ axis is defined as $1-R^{2}:=\frac{\mathbb{E}\|WX-\tilde{W}X\|^{2}}{\mathrm{Var}(WX)}$, where $X$ contains activation inputs from 101010 sequences of length 204820482048 from wikitext2. The right plot is bottom right corner of the left plot, zoomed in. We note that by paying a small price in the accuracy of the weight, we can reduce the amplification ratio dramatically.
Proof of Lemma 4.2.
Recalling the definition of $\tilde{W}=WH(H+J)^{-1}$, that $X$ and $Z$ are statistically independent, and $Z$ has zero mean, and that $H,J$ are positive semi-definite symmetric matrices, we have
$\mathbb{E}\|\delta(U)\|^{2}$ $=\mathbb{E}\|(W-U)X-UZ\|^{2}=(W-U)H(W-U)^{\top}+UJU^{\top}$ (7)
$=(\tilde{W}-U)(H+J)(\tilde{W}-U)+C(W,H,J),$ (8)
where
$C(W,H,J)=W(H-H(H+J)^{-1}H)W^{\top},$ (9)
is independent of $U$, and therefore does not affect the minimization. ∎
Appendix C Gosset oracle